# Primordial magnetogenesis in loop quantum cosmology

### v1

@@ I Introduction
It has been shown that electromagnetic fields of strengths observed in intergalactic voids can be generated in the early universe during inflation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] or other primordial scenarios [33, 34, 35, 36, 37] (for texts and reviews, see, for instance, [38, 39, 40, 41, 42, 43]). See also [44] in this context. In a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) universe, standard electromagnetic action is conformally invariant. Hence, the electromagnetic fields are not affected by the spacetime curvature of the universe. This implies that as the universe expands nearly exponentially during inflation, the energy density of magnetic fields will get diluted very quickly. This difficulty in generating magnetic fields can be overcome if we break the conformal invariance. There are various ways to break conformal invariance; one simple way is to couple the electromagnetic field to the background dynamics of the scalar field that drives the primordial universe, namely the inflaton (see, for instance, [17, 18, 21, 22]). This will ensure that magnetic fields present in the early universe will not get diluted during inflation. In this scenario, magnetic fields provide a new window to learn about the early universe through their coupling to the background.
Loop quantum cosmology (LQC) is an attempt to extend the inflationary phase to the Planck regime [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. For reviews on LQC, refer, for instance, [58, 59]. In LQC, inflation is preceded by a quantum bounce. The study of the magnetic field in LQC is interesting due to at least two reasons. First, it would be interesting to understand whether the magnetic field generated in LQC can lead to the observed magnetic fields. Second, magnetic fields generated in LQC could carry signatures of the quantum bounce. These signatures could provide vital information about the Planck regime. If tiny anisotropies are present in the background, they could grow during the contracting phase, leading to the bounce, and the spacetime during the quantum bounce could be anisotropic (see, for instance, [60, 61, 62, 63, 64, 65, 66]). There have been some efforts in understanding the generation of magnetic fields in this context (see, for instance, [67, 68, 69]). In this article, we shall take a different approach and study the generation of a magnetic field in LQC, assuming a homogeneous and isotropic spacetime. We shall consider the electromagnetic field as a test field coupled to the background dynamics in LQC.
@@
We shall work with natural units, i.e., $c\,=\,\hbar\,=\,1$. All masses will be expressed in Planck mass ${\rm m}_{\rm Pl}\,=\,1/\sqrt{G}$, where $G$ is Newton’s gravitational constant. Depending upon convenience, we shall work with cosmic time ($t$), conformal time ($\eta$), or e-fold ($N$). Different times are related to each other through ${\rm d}N\,=\,H(t)\,{\rm d}t\,=\,a(\eta)\,H(\eta)\,{\rm d}\eta$, where $a$ is the scale factor and $H$ is the Hubble parameter. For brevity, whenever there is no ambiguity, we shall omit explicitly expressing the time or spatial dependence. An overdot and prime will refer to a derivative with respect to cosmic time and conformal time, respectively. Throughout this article, we will use the suffixes ‘e’ and ‘0’ to denote a quantity computed at the end of inflation and today, respectively. For instance, $a_{e}$ and $a_{0}$ will refer to the value of the scale factor at the end of inflation and today, respectively. Greek indices such as $\mu,\,\nu,\dots$ imply either a time or spatial index. Latin indices such as $i,\,j,\,\dots$ will indicate only the spatial index. Finally, the magnetic field will be expressed in units of gauss (G).
@@ II Magnetogenesis in FLRW spacetime
The expectation value of the energy density of electric and magnetic field generated in this process can be found from the respective parts of the ‘$00$’ component of the energy-momentum tensor. The energy density of the magnetic field can be computed to be
@@ III Loop Quantum Cosmology
Loop quantum cosmology is an attempt based on principles of loop quantum gravity to extend the inflationary scenario into the Planck regime. In LQC, the inflationary phase is preceded by a quantum bounce. The quantum gravitational effects are effectively captured by the modified Friedmann - Raychaudhuri equations (see, for instance, [58, 59]), viz.,
@@
If the scalar field is governed by a suitable potential, then inflation can set in some time after the bounce [48, 71, 72, 73].
In this work, we will consider two potentials, namely, the quadratic potential [74] and the Starobinsky potential [75]. The former is chosen because of its simplicity, and the latter for being one of the best-known models of slow-roll inflation [76]. The quadratic potential is given by
@@
Loop quantum cosmology has another free parameter, namely the value of the scalar field at the bounce, denoted as $\phi_{b}$. For a given potential, the value of the scalar field at bounce determines the amount of expansion after the bounce till the time at which the pivot scale $k_{\star}\,=\,0.002{\rm Mpc}^{-1}$ leaves the horizon [53, 54]. More expansion prior to the exit of the pivot scale implies that signals from the bounce will not be visible at $k_{\star}$. In other words, more pre-inflationary expansion redshifts any signatures of the bounce away from the observable wavelengths. Pivot scale corresponds to a multipole of $\ell\approx 30$ in the cosmic microwave background. In the CMB power spectrum, most of the departure from near scale invariance occurs at multipoles of $\ell\lesssim 30$. We can put a lower bound on the value of $\phi_{b}$ by demanding that imprints of bounce are only visible at $k\lesssim k_{\star}$ (i.e., at multipoles $\ell\lesssim 30$). This corresponds, roughly, to setting $k_{\star}\,\approx 3\,k_{\rm LQC}$, where $k_{\rm LQC}\,=\,a(t_{B})\,\sqrt{8\,\pi\,\rho_{B}/{\rm m}_{\rm Pl}^{2}}\,\approx\,3.21{\rm m}_{\rm Pl}$. Following this approach, we have set $\phi_{b}\,=1.014{\rm m}_{\rm Pl}$ in the case of quadratic potential and $\phi_{b}\,=-1.4243\,{\rm m}_{\rm Pl}$ for Starobinsky potential. The velocity of the scalar field has been set to $\dot{\phi}_{b}\,=\,\sqrt{2\left(\rho_{\rm sup}\,-\,V(\phi_{b})\right)}$. Such an initial condition will lead to about $14.00$ e-folds of expansion from bounce till exit of pivot scale and $56.88$ e-folds of expansion during slow-roll for the case of quadratic potential. For the Starobinsky potential, this corresponds to a pre-inflationary expansion of about $15.86$ e-folds and $55.74$ e-folds since the exit of the pivot scale.
@@ IV Numerical evolution of magnetic fields in LQC
${\cal A}_{k}^{\prime\prime}(\eta)\,+\,\left(\,k^{2}\,-\,\frac{f^{\prime\prime}(\eta)}{f(\eta)}\right)\,=\,0.$ (14)
Note that this is an equation of motion of a time-dependent harmonic oscillator. The evolution of the mode, hence, depends on the form of the coupling function. If the coupling function f= 1f\,=\,1, then the mode just oscillates and doesn’t grow. Inspired from previous literature [21], we shall consider the following form of coupling function,
@@ V.1 Origin of scale dependence
The origin of the scale dependence of the vector potential can be understood by studying Eqn. (14). From this equation, we can see that, when $k^{2}>>f^{\prime\prime}/f$ the modes are simply oscillatory and when $k^{2}<<f^{\prime\prime}/f$, the modes evolve as ${\cal A}_{k}\,\propto\,f$. Thus, the evolution of the Fourier mode of the vector potential can be understood by comparing the term $\sqrt{|f^{\prime\prime}/f|}$ with the wavenumber of perturbations. We plot $\sqrt{|f^{\prime\prime}/f|}$ as a function of time, for both potentials of scalar fields in figure 2. We have also plotted comoving wavenumbers, as horizontal dashed and dotted lines, for comparison. Here, $k_{b}$ is a scale set by the peak of the function $\sqrt{|f^{\prime\prime}/f|}$ at the bounce. Note that $k_{b}\approx k_{LQC}$ for the coupling function we have chosen. If we impose Minkowski initial conditions at the bounce, we can see from the plot that modes $k\lesssim k_{b}$ are excited during the bounce and hence are not in the Minkowski vacuum at the onset of inflation. Whereas, modes $k>>k_{b}$ are not excited during the bounce and hence remain in the Minkowski vacuum at the onset of inflation. Thus, the behavior of modes with $k>>k_{b}$ will be similar to the behavior of these modes in slow-roll inflation. For the coupling function Eqn. (15), with $n=2$, we expect nearly scale-invariant magnetic power spectra for modes $k>>k_{b}$ and scale-dependent behaviour for smaller wavenumbers. From figure 1 we see that the magnetic spectrum of modes $k>>k_{b}\simeq k_{\star}/3$ is nearly scale-invariant, whereas intermediate modes with $k_{I}\lesssim k<k_{b}$ behave on average as $k^{-2}$. Here, $k_{I}$ is the wavenumber which becomes smaller than $\sqrt{|f^{\prime\prime}/f|}$, i.e., becomes ‘super-horizon’, just before the onset of inflation. Power spectra briefly behave as $1/k$ for wavenumbers around $k_{b}$, i.e., as the behaviour transitions from nearly scale invariant to a $k^{-2}$ behaviour. Wavenumbers with $k<<k_{I}$ satisfy the condition $k^{2}<<f^{\prime\prime}/f$ throughout its evolution and hence for these modes, ${\cal A}_{k}\,\propto\,f$. Thus, for $k<<k_{I}$, we see from Eqn. (7) that, for n= 2n\,=\,2, the magnetic spectra behaves as $k^{4}$. This behaviour of infrared modes is also evident from the figure 1.
@@ V.2 Effect of imposing initial conditions at different time
In obtaining figures 1 and 3 we imposed Minkowski initial condition at the bounce, i.e., at $t\,=\,0\,{\rm t}_{\rm Pl}$. In this section, we study the dependence of magnetic power spectra on the time at which Minkowski initial conditions are imposed. Figures 4 and 5 contain graphs of magnetic power spectra generated in LQC with quadratic and Starobinsky models, respectively, upon imposing Minkowski initial conditions at different times. The spectra in each case have been computed after modes become $k<<\sqrt{|f^{\prime\prime}/f|}$, i.e., ‘super-horizon’, during inflation. Note that this is equivalent to computing the magnetic spectrum at the end of inflation. This is because the magnetic power spectrum remains constant after the mode becomes ‘super-horizon’ for the case of the coupling function with $n=2$.
Figure 4 corresponds to the magnetic spectra generated in the case of the quadratic potential. We find that when initial conditions are imposed before the bounce, the obtained spectra are slightly higher than the one obtained upon imposing initial conditions at the bounce. However, we find that the spectra obtained with initial conditions imposed at any time prior to the bounce are largely similar. This may be because of the sharp drop in the amplitude of $\sqrt{|f^{\prime\prime}/f|}$ before the bounce (see figure 2). When initial conditions are imposed after the bounce, we see that the scale dependence decreases a lot and the spectra become similar to that obtained during inflation. This can again be attributed to the sharp drop in the amplitude of $\sqrt{|f^{\prime\prime}/f|}$ after the bounce. The scale dependence of infrared modes, i.e., $k\,<\,k_{I}$, remains the same regardless of the time at which the initial condition is imposed. As expected, they display a $k^{4}$ dependence since, as explained in the previous section, these modes are at most of the time, ‘super-horizon’.
@@
Figure 5 plots magnetic power spectra generated in the Starobinsky model when the initial conditions are imposed at different times. As in the case of the quadratic potential, we impose the Minkowski initial condition, and the spectra are computed when modes are ‘super-horizon’. The dependence of power spectra generated in the case of the Starobinsky model is more involved. This is expected and can be understood from the behaviour of $\sqrt{|f^{\prime\prime}/f|}$ given in figure 2. As in the case of quadratic potential, the spectra generated when initial conditions are imposed before the bounce are larger than that obtained when initial condition is imposed at the bounce. However, unlike in the case of quadratic potential, the amplitude of the spectrum for modes $k_{I}<k<<k_{b}$ depends strongly on the time at which initial conditions are imposed. We find that the earlier the initial time, the larger the amplitude of the spectrum for these scales. This behaviour may be attributed to the presence of an additional bump in $\sqrt{|f^{\prime\prime}/f|}$, see figure 2, for the Starobinsky model. The extra bump excites the modes even before they are excited during the bounce. The behaviour of $\sqrt{|f^{\prime\prime}/f|}$ after the bounce is similar in both potentials. Hence, the dependence of the power spectrum on the initial time, if the initial conditions are imposed after the bounce, is similar to that obtained in the case of the quadratic potential. The scale dependence of the power spectrum decreases a lot and becomes largely similar to that obtained in the context of slow-roll inflation if the initial time is farther from the bounce. As in the case of the quadratic potential, the spectra of infrared modes behave as $k^{4}$ irrespective of the time at which the initial condition is imposed.

### v2

@@ I Introduction
It has been shown that electromagnetic fields of strengths observed in intergalactic voids can be generated in the early Universe during inflation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] or other primordial scenarios [32, 33, 34, 35, 36, 37, 38] (for texts and reviews, see, for instance, [39, 40, 41, 14, 42, 43]). See also [44, 45] in this context. In a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) universe, standard electromagnetic action is conformally invariant. Hence, the electromagnetic fields are not affected by the spacetime curvature of the universe. This implies that as the universe expands nearly exponentially during inflation, the energy density of magnetic fields will get diluted very quickly. This difficulty in generating magnetic fields can be overcome if we break the conformal invariance. There are various ways to break conformal invariance; one simple way is to couple the electromagnetic field to the background dynamics of the scalar field that drives the primordial universe, namely the inflaton (see, for instance, [17, 18, 22, 23]). This will ensure that magnetic fields present in the early Universe will not get diluted during inflation. In this scenario, magnetic fields provide a new window to learn about the early Universe through their coupling to the background.
Loop quantum cosmology (LQC) is an attempt to extend the inflationary phase to the Planck regime [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. For reviews on LQC, refer, for instance, [62, 63]. In LQC, inflation is preceded by a quantum bounce. The study of the magnetic field in LQC is interesting due to at least two reasons. First, it would be interesting to understand whether the magnetic field generated in LQC can lead to the observed magnetic fields. Second, magnetic fields generated in LQC could carry signatures of the quantum bounce. These signatures could provide vital information about the Planck regime. If tiny anisotropies are present in the background, they could grow during the contracting phase, leading to the bounce, and the spacetime during the quantum bounce could be anisotropic (see, for instance, [64, 65, 66, 67, 68, 69, 70]). There have been efforts to understand the generation of homogeneous magnetic fields in this context (see, for instance, [71, 72, 73]). In this article, we shall take a different approach and study the generation of an inhomogeneous magnetic field in LQC, assuming a homogeneous and isotropic spacetime. We shall consider the electromagnetic field as a test field coupled to the background dynamics in LQC.
@@
We shall work with natural units, i.e., $c\,=\,\hbar\,=\,1$. All masses will be expressed in Planck mass ${\rm m}_{\rm Pl}\,=\,1/\sqrt{G}$, where $G$ is Newton’s gravitational constant. Depending upon convenience, we shall work with cosmic time ($t$), conformal time ($\eta$), or e-fold ($N$). Different times are related to each other through ${\rm d}N\,=\,H(t)\,{\rm d}t\,=\,a(\eta)\,H(\eta)\,{\rm d}\eta$, where $a$ is the scale factor and $H$ is the Hubble parameter. For brevity, whenever there is no ambiguity, we shall omit explicitly expressing the time or spatial dependence. An overdot and prime will refer to a derivative with respect to cosmic time and conformal time, respectively. Throughout this article, we will use the suffixes ‘e and 0’ to denote a quantity computed at the end of inflation and today, respectively. For instance, $a_{e}$ and $a_{0}$ will refer to the value of the scale factor at the end of inflation and today, respectively. Greek indices such as $\mu,\,\nu,\dots$ imply either a time or spatial index. Latin indices such as $i,\,j,\,\dots$ will indicate only the spatial index. Finally, the magnetic field will be expressed in units of gauss (G).
@@ II Magnetogenesis in FLRW spacetime
The expectation value of the energy density of the electric and magnetic fields generated in this process can be found from the respective parts of the ‘$00$’ component of the energy-momentum tensor. The energy density of the magnetic field can be computed to be
@@ III Loop Quantum Cosmology
Loop quantum cosmology is an attempt, based on the principles of loop quantum gravity, to extend the inflationary scenario into the Planck regime (for reviews, see, for instance, [62, 63]). In LQC, a FLRW spacetime sourced by a scalar field is defined by a quantum wave function, which depends only on the scalar field and the volume, where volume is proportional to the cube of the scale factor. We will consider the case where energy density in the quantum regime is dominated by the kinetic term of the scalar field [47, 51, 62]. As we will see, such kinetic-dominated bounces could lead to sufficient e-folds of inflation after the bounce [74, 75, 76]. Among the class of solutions admitted by LQC, we are interested in states that are sharply peaked around the classical trajectories, particularly those with small dispersion in volume, at late times. It has been shown that [48, 49, 55, 56, 77], for such states, the quantum dynamics are effectively captured by the modified Friedmann-Raychaudhuri equations, viz.,
@@
Subsequent evolution after the bounce quickly decreases the energy density and makes the terms in Eqns. (10) which are nonlinear in energy density, irrelevant, and these equations will reduce to the familiar Friedmann-Raychaudhuri equations in general relativity. Hence, if the scalar field is governed by a suitable potential, then slow roll inflation can set in some time after the bounce [50, 74, 75, 76].
In this work, we will consider two potentials for the scalar field, namely, the quadratic potential [79] and the Starobinsky potential [80]. The former is chosen because of its simplicity, and the latter for being one of the best-known models of slow-roll inflation [81]. The quadratic potential is given by
@@
Refer to caption Refer to caption
Figure 1: Behaviour of scale factor (top), scalar field (bottom left), and energy density of scalar field (bottom right) in LQC for the initial conditions described in the text for the two potentials. From the plot of $a$, note that the epoch of inflation is preceded by a bounce, where an expanding phase transitions to a contracting phase. From the plot of $\phi$ we see that the scalar field reaches a maximum value sometime after the bounce and then decreases with time. For the scenario considered here, inflation is achieved during this epoch. From the plot of $\rho$, we see that energy density is maximum at the bounce. Further, we see from the inset that the behaviour of the scale factor near the bounce is independent of the form of the potential.
Loop quantum cosmology has another free parameter, namely the value of the scalar field at the bounce, denoted as $\phi_{b}$. For a given potential, the value of the scalar field at bounce determines the amount of expansion after the bounce till the time at which the pivot scale $k_{\star}\,=\,0.002{\rm Mpc}^{-1}$ leaves the horizon [57, 58]. More expansion prior to the exit of the pivot scale implies that signals from the bounce will not be visible at $k_{\star}$. In other words, more preinflationary expansion redshifts any signatures of the bounce away from the observable wavelengths. Pivot scale corresponds to a multipole of $\ell\approx 30$ in the cosmic microwave background. In the CMB power spectrum, most of the departure from near scale invariance occurs at multipoles of $\ell\lesssim 30$. We can put a lower bound on the value of $\phi_{b}$ by demanding that imprints of bounce are only visible at $k\lesssim k_{\star}$ (i.e., at multipoles $\ell\lesssim 30$). This corresponds, roughly, to setting $k_{\star}\,\approx 3\,k_{\rm LQC}$, where $k_{\rm LQC}\,=\,a(t_{B})\,\sqrt{8\,\pi\,\rho_{B}/{\rm m}_{\rm Pl}^{2}}\,\approx\,3.21{\rm m}_{\rm Pl}$. Following this approach, we have set $\phi_{b}\,=1.014{\rm m}_{\rm Pl}$ in the case of quadratic potential and $\phi_{b}\,=-1.4243\,{\rm m}_{\rm Pl}$ for Starobinsky potential. The velocity of the scalar field has been set to $\dot{\phi}_{b}\,=\,\sqrt{2\left(\rho_{\rm sup}\,-\,V(\phi_{b})\right)}$. Such an initial condition will lead to about $14.00$ e-folds of expansion from bounce till exit of pivot scale and $56.88$ e-folds of expansion during slow roll for the case of quadratic potential. For the Starobinsky potential, this corresponds to a preinflationary expansion of about $15.86$ e-folds and $55.74$ e-folds since the exit of the pivot scale.
The dynamics of background quantities $a$, $\phi$, and energy density $\rho$ in LQC for the two potentials with the initial conditions mentioned above are given in figure 1. The figure illustrates that in LQC, the scale factor decreases initially, reaches a minimum value, undergoes a bounce, and then experiences inflationary expansion. From the plot of $\phi$, we see that, for the initial conditions we have considered, after the bounce, the kinetic energy of the scalar field is spent in climbing up the potential. After a while field reaches a maximum and starts rolling down. Slow-roll inflation sets in as the field begins to roll down. Inflation ends as the inflaton approaches the bottom of its potential. The scalar field then undergoes damped oscillations about that minimum. The figure also illustrates the behaviour of the energy density of the scalar field as a function of time. From the plots of $a$ and $\rho$, it is clear that energy density is maximum and equal to $\rho_{\rm sup}$ at the bounce. The energy density then quickly decays, and the quantum effects become negligible. The nearly flat behavior of the energy density indicates the epoch during which inflation occurs. Insets of scale factor and the energy density of the scalar field show that their behaviour is independent of the potential near the bounce.
@@ IV Numerical evolution of magnetic fields in LQC
${\cal A}_{k}^{\prime\prime}(\eta)\,+\,\left(\,k^{2}\,-\,\frac{f^{\prime\prime}(\eta)}{f(\eta)}\right){\cal A}_{k}(\eta)\,=\,0.$ (14)
Note that this is an equation of motion of a time-dependent harmonic oscillator. The evolution of the mode, hence, depends on the form of the coupling function. If the coupling function f= 1f\,=\,1, then the mode just oscillates and does not grow. Inspired from previous literature [22], we shall consider the following form of coupling function,
@@ V.1 Origin of scale dependence
The origin of the scale dependence of the vector potential can be understood by studying Eqn. (14). From this equation, we can see that, when $k^{2}>>f^{\prime\prime}/f$ the modes are simply oscillatory and when $k^{2}<<f^{\prime\prime}/f$, the modes evolve as ${\cal A}_{k}\,\propto\,f$. Thus, the evolution of the Fourier mode of the vector potential can be understood by comparing the term $\sqrt{|f^{\prime\prime}/f|}$ with the wave number of perturbations. We plot $\sqrt{|f^{\prime\prime}/f|}$ as a function of time, for both potentials of scalar fields in figure 3. We have also plotted comoving wave numbers, as horizontal dashed and dotted lines, for comparison. Here, $k_{b}$ is a scale set by the peak of the function $\sqrt{|f^{\prime\prime}/f|}$ at the bounce. Note that $k_{b}\approx k_{LQC}$ for the coupling function we have chosen. If we impose Minkowski initial conditions at the bounce, we can see from the plot that modes $k\lesssim k_{b}$ are excited during the bounce and hence are not in the Minkowski vacuum at the onset of inflation. Whereas, modes $k>>k_{b}$ are not excited during the bounce and hence remain in the Minkowski vacuum at the onset of inflation. Thus, the behavior of modes with $k>>k_{b}$ will be similar to the behavior of these modes in slow-roll inflation. For the coupling function Eqn. (15), with $n=2$, we expect nearly scale-invariant magnetic power spectra for modes $k>>k_{b}$ and scale-dependent behaviour for smaller wave numbers. From figure 2 we see that the magnetic spectrum of modes $k>>k_{b}\simeq k_{\star}/3$ is nearly scale-invariant, whereas intermediate modes with $k_{I}\lesssim k<k_{b}$ behave on average as $k^{-2}$. Here, $k_{I}$ is the wave number which becomes smaller than $\sqrt{|f^{\prime\prime}/f|}$, i.e., becomes superhorizon, just before the onset of inflation. Power spectra briefly behave as $1/k$ for wave numbers around $k_{b}$, i.e., as the behaviour transitions from nearly scale invariant to a $k^{-2}$ behaviour. wave numbers with $k<<k_{I}$ satisfy the condition $k^{2}<<f^{\prime\prime}/f$ throughout its evolution and hence for these modes, ${\cal A}_{k}\,\propto\,f$. Thus, for $k<<k_{I}$, we see from Eqn. (7) that, for n= 2n\,=\,2, the magnetic spectra behaves as $k^{4}$. This behaviour of infrared modes is also evident from the figure 2.
@@ V.2 Effect of imposing initial conditions at different time
In obtaining figures 2 and 4 we imposed Minkowski initial condition at the bounce, i.e., at $t\,=\,0\,{\rm t}_{\rm Pl}$. In this section, we study the dependence of magnetic power spectra on the time at which Minkowski initial conditions are imposed. Figures 5 and 6 contain graphs of magnetic power spectra generated in LQC with quadratic and Starobinsky models, respectively, upon imposing Minkowski initial conditions at different times. The spectra in each case have been computed after modes become $k<<\sqrt{|f^{\prime\prime}/f|}$, i.e., superhorizon, during inflation. Note that this is equivalent to computing the magnetic spectrum at the end of inflation. This is because the magnetic power spectrum remains constant after the mode becomes superhorizon for the case of the coupling function with $n=2$.
Figure 5 corresponds to the magnetic spectra generated in the case of the quadratic potential. We find that when initial conditions are imposed before the bounce, the obtained spectra are slightly higher than the one obtained upon imposing initial conditions at the bounce. However, we find that the spectra obtained with initial conditions imposed at any time prior to the bounce are largely similar. This may be because of the sharp drop in the amplitude of $\sqrt{|f^{\prime\prime}/f|}$ before the bounce (see figure 3). When initial conditions are imposed after the bounce, we see that the scale dependence decreases a lot and the spectra become similar to that obtained during inflation. This can again be attributed to the sharp drop in the amplitude of $\sqrt{|f^{\prime\prime}/f|}$ after the bounce. The scale dependence of infrared modes, i.e., $k\,<\,k_{I}$, remains the same regardless of the time at which the initial condition is imposed. As expected, they display a $k^{4}$ dependence since, as explained in the previous section, these modes are at most of the time, superhorizon.
@@
Figure 6 plots magnetic power spectra generated in the Starobinsky model when the initial conditions are imposed at different times. As in the case of the quadratic potential, we impose the Minkowski initial condition, and the spectra are computed when modes are superhorizon. The dependence of power spectra generated in the case of the Starobinsky model is more involved. This is expected and can be understood from the behaviour of $\sqrt{|f^{\prime\prime}/f|}$ given in figure 3. As in the case of quadratic potential, the spectra generated when initial conditions are imposed before the bounce are larger than that obtained when initial condition is imposed at the bounce. However, unlike in the case of quadratic potential, the amplitude of the spectrum for modes $k_{I}<k<<k_{b}$ depends strongly on the time at which initial conditions are imposed. We find that the earlier the initial time, the larger the amplitude of the spectrum for these scales. This behaviour may be attributed to the presence of an additional bump in $\sqrt{|f^{\prime\prime}/f|}$, see figure 3, for the Starobinsky model. The extra bump excites the modes even before they are excited during the bounce. The behaviour of $\sqrt{|f^{\prime\prime}/f|}$ after the bounce is similar in both potentials. Hence, the dependence of the power spectrum on the initial time, if the initial conditions are imposed after the bounce, is similar to that obtained in the case of the quadratic potential. The scale dependence of the power spectrum decreases a lot and becomes largely similar to that obtained in the context of slow-roll inflation if the initial time is farther from the bounce. As in the case of the quadratic potential, the spectra of infrared modes behave as $k^{4}$ irrespective of the time at which the initial condition is imposed.